Circular No. 6078

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

X-RAY NOVA IN SCORPIUS J. Greiner, Max Planck Institute for Extraterrestrial Physics, reports:

“The x-ray transient GRO J1655–40 was observed with the ROSAT HRI for a total of 785 s, starting at Aug. 28.6616 UT. The best-fit x-ray position is determined to be α = 16^{h}53^{m}59^{s}.9, δ = −39^{o}50′44′′ (equinox 2000.0), with a (mainly systematic) error of ± 8′′. This is only 2′′ off the radio (IAUC 6055) and optical (IAUC 6060) positions. The mean x-ray flux is rather constant at a HRI count rate of ∼ 65 counts/s (corresponding to ∼ 300 mCrab in the range 0.1–2.4 keV) during the observation interval, consistent with ASCA observations on Aug. 23 (IAUC 6063). Superimposed on this mean level are x-ray flux variations between 20 and 150 counts/s on timescales down to 0.1 s. Previous ROSAT PSPC observations did not show this source; the upper limits are 0.02 count/s (Aug. 1990, All-Sky-Survey) and 0.003 count/s (Sept. 1993, courtesy U. Herbstmeier). Adopting the absorption column density of 5 × 10^{21} cm^{−2} and a power-law photon index of −2.1, as determined with ASCA (IAUC 6063), the latter upper limit corresponds to an unabsorbed flux of 2.7 × 10^{−13} erg cm^{−2} s^{−1} in the range 0.1–2.4 keV — a factor of 2.3 × 10^{4} lower than the mean flux during the HRI observation.”

D. Campbell-Wilson, University of Sydney and Molonglo Observatory Synthesis Telescope (MOST); D. J. McKay, Australia Telescope National Facility; and J. E. Lovell, University of Tasmania, report on continued flux-density monitoring of the radio source associated with GRO J1655–40: “The source is undergoing a second major radio outburst. Observations with the MOST and the AT Compact Array revealed the presence of an optically-thick, flat-spectrum radio source that appeared on Sept. 9 (see Hjellming and Rupen, IAUC 6077). This new source had a radio spectrum that peaked at 0.45 Jy at 5 GHz on Sept. 11. On Sept. 14.32 UT, the source reached 1.6 Jy at 843 MHz, rising to 2.1 Jy on Sept. 14.45. Observations at Mt. Pleasant Observatory at 2.3, 6.7, and 8.4 GHz show the spectrum of the new outburst to be optically thin. This second outburst follows the increased x-ray emission seen by the BATSE and GRANAT teams (IAUC 6075, 6077). The second outburst suggests a binary origin for GRO J1655– 40, with a period of 31 ± 1 days, based on the timing of the previous radio flare (Campbell-Wilson and Hunstead, IAUC 6055).”

1994 September 15 Daniel W. E. Green